OpenStudy (anonymous):

To how many decimal places is the symmetric difference quotient accurate when it is used to approximate f ‘(0) for f(x) =4^x and h=0.08?

OpenStudy (kc_kennylau):

\[f'(x)\approx\frac{\Delta x}{\Delta h}\]

OpenStudy (kc_kennylau):

\[=\frac{f(x+h)-f(x)}h\]\[\large=\frac{4^{x+h}-4^x}h\]

OpenStudy (kc_kennylau):

now plug the values in :)

OpenStudy (anonymous):

ok i got 2?

OpenStudy (anonymous):

@kc_kennylau